OpenStudy (anonymous):

the rectangular prism below measures 4in by 6in by X. it the surface area is 308 square inches what is the hight if the prism

OpenStudy (anonymous):

Surface Area of Rec. Prism = 2(LW+LH+WH) 308=2(4x6+4xh+6xh) 308=2(24+10h) 154=24+10h 10h=154-24 h=130/10 h=13